# Multi-Modal Data-Efficient 3D Scene Understanding for Autonomous Driving

### v2

@@ Abstract
Efficient data utilization is crucial for advancing 3D scene understanding in autonomous driving, where reliance on heavily human-annotated LiDAR point clouds challenges fully supervised methods. Addressing this, our study extends into semi-supervised learning for LiDAR semantic segmentation, leveraging the intrinsic spatial priors of driving scenes and multi-sensor complements to augment the efficacy of unlabeled datasets. We introduce LaserMix++, an evolved framework that integrates laser beam manipulations from disparate LiDAR scans and incorporates LiDAR-camera correspondences to further assist data-efficient learning. Our framework is tailored to enhance 3D scene consistency regularization by incorporating multi-modality, including 1) multi-modal LaserMix operation for fine-grained cross-sensor interactions; 2) camera-to-LiDAR feature distillation that enhances LiDAR feature learning; and 3) language-driven knowledge guidance generating auxiliary supervisions using open-vocabulary models. The versatility of LaserMix++ enables applications across LiDAR representations, establishing it as a universally applicable solution. Our framework is rigorously validated through theoretical analysis and extensive experiments on popular driving perception datasets. Results demonstrate that LaserMix++ markedly outperforms fully supervised alternatives, achieving comparable accuracy with five times fewer annotations and significantly improving the supervised-only baselines. This substantial advancement underscores the potential of semi-supervised approaches in reducing the reliance on extensive labeled data in LiDAR-based 3D scene understanding systems. Code and benchmark toolkits are publicly available at https://github.com/ldkong1205/LaserMix.
Index Terms:
Semi-Supervised Learning; LiDAR Semantic Segmentation; 3D Scene Understanding; Autonomous Driving; Robustness
@@ 1 Introduction
Figure 1: Motivation. (a) We observe a strong spatial prior from LiDAR-acquired driving scenes, where objects and backgrounds around the ego-vehicle have a patterned distribution on different (lower, middle, and upper) laser beams. (b) The proposed laser beam mixing technique is agnostic to different LiDAR modalities and can be universally applied to existing LiDAR segmentation backbones. (c) Our approaches achieved superior performance than state-of-the-art methods [33, 34, 35, 36, 37] under both low-data (10%, 20%, and 50% labels) and high-data (full supervision) regimes on nuScenes [38].
@@ 3 Data-Efficient 3D Scene Understanding
In this section, we first introduce the spatial prior of LiDAR-based driving scenes (Sec. 3.1). We then present LaserMix, which strives to efficiently encourage confident and consistent LiDAR predictions (Sec. 3.2). Finally, we establish a strong 3D scene consistency regularization baseline (Sec. 3.3).
@@ 3.1 Spatial Prior in 3D Scene Understanding
Spatial Prior Observation. Understanding and utilizing the spatial distribution inherent in LiDAR scenes is pivotal for enhancing semi-supervised learning. LiDAR point clouds offer unique spatial priors that are not prevalent in 2D images. As shown in Tab. I, each LiDAR-acquired semantic class poses unique distribution patterns that directly reflect real-world driving scenes. Our methodology leverages these priors by encouraging the model to maintain consistent predictions across varying LiDAR data manipulations.
TABLE I: A case study on the strong spatial prior of representative semantic classes from the SemanticKITTI [28] dataset. For each class, we show its type (static or dynamic), proportion (valid number of points in percentage), distribution among eight areas ($A=\{a_{1},a_{2},...,a_{8}\}$, i.e., eight laser beam groups), and the heatmap in range view (lighter colors correspond to areas that have a higher likelihood to appear and vice versa). Best viewed in colors and zoom-ed in for additional details.
@@
motorcyclist dynamic $0.004\%$ [Uncaptioned image] [Uncaptioned image]
@@
where $c$ is a small constant and $\theta$ represents the model parameters. Similar to the classic entropy minimization framework [86], the constraint in Equation 1 is transformed into a probabilistic model where the model parameter distribution is guided by the principle of maximum entropy:
$\begin{split}P(\theta)&\propto\exp(-\lambda H(X_{\mathrm{in}},Y_{\mathrm{in}}|A))\\ &\propto\exp(-\lambda H(Y_{\mathrm{in}}|X_{\mathrm{in}},A))\,,\end{split}$ (2)
where $\lambda>0$ acts as the Lagrange multiplier associated with the entropy constraint, which corresponds to constant $c$. $H(X_{\mathrm{in}}|A)$ has been ignored for being independent of the model parameter $\theta$. Here, we consider Equation 2 as the formal formulation of the spatial prior and discuss how to empirically compute it in the following sections.
Marginalization. To utilize the spatial prior defined in Equation 2, we empirically compute the entropy $H(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ of the LiDAR points inside area $A$ as follows:
$\begin{split}&\hat{H}(Y_{\mathrm{in}}|X_{\mathrm{in}},A)=\\ &\hat{\mathbb{E}}_{X_{\mathrm{in}},Y_{\mathrm{in}},A}[P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)\log P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)]\,,\end{split}$ (3)
where $\hat{.}$ denotes the empirical estimation. The end-to-end LiDAR segmentation model $\mathcal{G}_{\theta}$ (with parameters $\theta$) usually takes full-sized data as inputs during the inference. Therefore, to compute $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ in LABEL:eq:empirical-entropy, we first pad the data outside the area to obtain the full-sized data. Here, we denote the data outside the area as $X_{\mathrm{out}}$; we then let the model infer $P(Y_{\mathrm{in}}|X_{\mathrm{in}},X_{\mathrm{out}},A)$, and finally marginalize $X_{\mathrm{out}}$ as:
@@
Training Objectives. Finally, we train the segmentation model $\mathcal{G}_{\theta}$ using the standard maximum-a-posteriori (MAP) estimation. We maximize the posterior that can be computed by Equation 2, LABEL:eq:empirical-entropy, and Equation 4, which is formulated as follows:
@@
Here, $L(\theta)$ is the likelihood function which can be computed using labeled data, i.e., the conventional supervised learning. Minimizing $\hat{H}(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ requires the marginal probability $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ to be confident, which further requires $P(Y_{\mathrm{in}}|X_{\mathrm{in}},X_{\mathrm{out}},A)$ to be both confident and consistent with respect to different outside data $X_{\mathrm{out}}$. To sum up, our proposed semi-supervised learning framework in Equation 5 encourages the segmentation model to make confident and consistent predictions at a predefined area, regardless of the data outside the area. The predefined area set $A$ determines the “strength” of the prior. When setting $A$ to the full area (i.e., the whole point cloud), our framework degrades to the classic entropy minimization framework [86].
@@ 3.2 LaserMix
where $(p^{x},p^{y},p^{z})$ represent the Cartesian coordinates of the LiDAR points. For any two LiDAR scans, $x_{1}$ and $x_{2}$, we first group all points from each scan by their inclination angles. More concretely, to establish $m$ non-overlapping areas, a set of $m+1$ inclination angles $\Phi=\{\phi_{0},\phi_{1},\phi_{2},...,\phi_{m}\}$ will be evenly sampled within the range of the minimum and maximum inclination angles in the dataset, and the area set $A=\{a_{1},a_{2},...,a_{m}\}$ can then be confined by bounding area $a_{i}$ in the inclination range $[\phi_{i-1},\phi_{i})$. It is important to note that the range of inclination angles varies depending on the specific configurations of different LiDAR sensors.
Role in our framework: The proposed laser-based partitioning aims to effectively “excite” a strong spatial prior in the LiDAR point cloud, as described by Step 1 in our semi-supervised learning framework. As shown in Figure 1(a) and Tab. I, this partitioning reveals clear distribution patterns in the semantic classes detected by each laser beam. Despite being an empirical choice, we will show in later sections that our laser-based partitioning method significantly outperforms other partition choices, including random points (MixUp-like partition [87]), random areas (CutMix-like partition [88]), and other heuristics like azimuth $\alpha$ (sensor horizontal direction) or radius $r$ (sensor range direction) partitions.
@@
$\begin{split}\tilde{x}_{1},\tilde{x}_{2}=\textrm{LaserMix}(x_{1},x_{2})\,,\\ \tilde{x}_{1}=x_{1}^{a_{1}}\cup x_{2}^{a_{2}}\cup x_{1}^{a_{3}}\cup\cdots,\\ \tilde{x}_{2}=x_{2}^{a_{1}}\cup x_{1}^{a_{2}}\cup x_{2}^{a_{3}}\cup\cdots,\\ \end{split}$ (7)
where $x_{i}^{a_{j}}$ is the data crop of $x_{i}$ confined within area $a_{j}$. Correspondingly, the semantic labels are mixed in the same way as Equation 7. It is worth highlighting that LaserMix is directly applied to the LiDAR point clouds and is thus agnostic to the various LiDAR representations, such as range view [40], bird’s eye view [37], raw points [89], sparse voxel [33], and multi-view fusion [41] representations. This versatility allows LaserMix to be effectively integrated across a broad spectrum of existing LiDAR segmentation frameworks, without necessitating any modifications to the underlying data or network structures. While LaserMix partitions the LiDAR point cloud for enhancing feature learning, concerns may arise regarding the integrity of object instances, such as cars, when these are split across different segments. However, given the large spatial area typically involved, the likelihood of significant object partitioning is low. Moreover, LaserMix’s use of both ground truth and high-confidence pseudo labels provides robust supervision, ensuring accurate object interpretation even when partitioning occurs. This design allows LaserMix to leverage the benefits of data augmentation without compromising object integrity, thereby enhancing the robustness and generalization of the model.
@@
Role in our framework: Central to our semi-supervised learning framework, LaserMix streamlines the computation of marginal probability $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ – a process otherwise computationally intensive in typical scenarios – as described by Step 2 in our framework. The cost for directly computing the marginal probability in Equation 4 on real-world LiDAR data is prohibitive; we need to iterate through all areas in $A$ and all outside data in $X_{\mathrm{out}}$, which requires $|A|\cdot|X_{\mathrm{out}}|$ predictions in total. To reduce the training overhead, we take advantage of the fact that a prediction in an area will be largely affected by its neighboring areas and let $X_{\mathrm{out}}$ fill only the neighbors instead of all the remaining areas. LaserMix mixes two scans by intertwining the areas so that the neighbors of each area are filled with data from the other scan. As a result, we obtain the prediction on all areas $A$ of two scans from only two predictions, which on average reduces the cost from $|A|$ to $1$. The scan before and after mixing counts as two data fillings, therefore $|X_{\mathrm{out}}|=2$. Overall, the training overhead is reduced from $|A|\cdot|X_{\mathrm{out}}|$ to $2$: only one prediction on original data and one additional prediction on mixed data are required for each LiDAR scan. During training, the memory consumption for a batch will be 2×2$\times$2 × compared to a standard semi-supervised learning framework, and the training speed will not be affected.
Figure 3: Overview of our baseline consistency regularization framework. We feed the labeled scan $x_{l}$ into the Student network to compute the supervised loss $\mathcal{L}_{\text{sup}}$ (w/ ground truth $y_{l}$). The unlabeled scan $x_{u}$ and the generated pseudo-label $y_{u}$ are mixed with $(x_{l},y_{l})$ via LaserMix (Sec. 3.2) to produce mixed data sample $(x_{\text{mix}},y_{\text{mix}})$, which is then fed into the Student network to compute the mixing loss $\mathcal{L}_{\text{mix}}$. Additionally, we encourage the consistency between the Student network and the Teacher network by computing the mean teacher loss $\mathcal{L}_{\text{mt}}$ over their predictions, where the Teacher network’s parameters are updated by the exponential moving average of that of the Student network. During inference, only the Teacher network is needed which maintains the same computational cost as the conventional LiDAR segmentation pipeline.
@@
$\mathcal{L}_{\text{mt}}=\frac{1}{|x|}\sum_{i=1}^{|x|}||\mathcal{G}_{\theta}^{s,(i)}(x)-\mathcal{G}_{\theta}^{t,(i)}(x)||^{2}_{2}\,,$ (9)
@@
Role in our framework: Our consistency regularization pipeline is designed to effectively minimize the marginal entropy for data-efficient 3D scene understanding, as described in Step 3 of our framework. Since the objective for minimizing the entropy has a challenging optimization landscape, pseudo-labeling emerges as a resort in practice [90]. Unlike conventional pseudo-label optimization in semi-supervised learning which only aims to encourage the predictions to be confident, minimizing the marginal entropy requires all predictions to be both confident and consistent. To this end, we leverage both ground truth and pseudo-labels as anchors to encourage the 3D scene understanding model’s predictions to be confident and consistent with these supervision signals.

### v3

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
Efficient data utilization is crucial for advancing 3D scene understanding in autonomous driving, where reliance on heavily human-annotated LiDAR point clouds challenges fully supervised methods. Addressing this, our study extends into semi-supervised learning for LiDAR semantic segmentation, leveraging the intrinsic spatial priors of driving scenes and multi-sensor complements to augment the efficacy of unlabeled datasets. We introduce LaserMix++, an evolved framework that integrates laser beam manipulations from disparate LiDAR scans and incorporates LiDAR-camera correspondences to further assist data-efficient learning. Our framework is tailored to enhance 3D scene consistency regularization by incorporating multi-modality, including 1) multi-modal LaserMix operation for fine-grained cross-sensor interactions; 2) camera-to-LiDAR feature distillation that enhances LiDAR feature learning; and 3) language-driven knowledge guidance generating auxiliary supervisions using open-vocabulary models. The versatility of LaserMix++ enables applications across LiDAR representations, establishing it as a universally applicable solution. Our framework is rigorously validated through theoretical analysis and extensive experiments on popular driving perception datasets. Results demonstrate that LaserMix++ markedly outperforms fully supervised alternatives, achieving comparable accuracy with five times fewer annotations and significantly improving the supervised-only baselines. This substantial advancement underscores the potential of semi-supervised approaches in reducing the reliance on extensive labeled data in LiDAR-based 3D scene understanding systems.
[…]
[ [Uncaptioned image] Project Page]https://ldkong.com/LaserMix
[…]
@@ 1 Introduction
Figure 1: Motivation. (a) We observe a strong spatial prior from LiDAR-acquired driving scenes, where objects and backgrounds around the ego-vehicle have a patterned distribution on different (lower, middle, upper) laser beams. (b) The proposed laser beam mixing technique is agnostic to different LiDAR modalities and can be universally applied to existing LiDAR segmentation backbones. (c) Our approaches achieved superior performance than state-of-the-art methods [107, 6, 15, 104, 102] under low-data (10%, 20%, 50% labels) and high-data (full supervision) regimes on nuScenes [18].
@@ 3 Data-Efficient 3D Scene Understanding
In this section, we first introduce the spatial prior of LiDAR-based driving scenes (Section 3.1). We then present LaserMix, which strives to efficiently encourage confident and consistent LiDAR predictions (Section 3.2). Finally, we establish a strong 3D scene consistency regularization baseline (Section 3.3).
@@ 3.1 Spatial Prior in 3D Scene Understanding
Spatial Prior Observation. Understanding and utilizing the spatial distribution inherent in LiDAR scenes is pivotal for enhancing semi-supervised learning. LiDAR point clouds offer unique spatial priors that are not prevalent in 2D images. As shown in Table 1, each LiDAR-acquired semantic class poses unique distribution patterns that directly reflect real-world driving scenes. Our methodology leverages these priors by encouraging the model to maintain consistent predictions across varying LiDAR data manipulations.
Table 1: A case study on the strong spatial prior of representative semantic classes from the SemanticKITTI [2] dataset. For each class, we show its type (static or dynamic), proportion (valid number of points in percentage), distribution among eight areas ($A=\{a_{1},a_{2},...,a_{8}\}$, i.e., eight laser beam groups), and the heatmap in range view (lighter colors correspond to areas that have a higher likelihood to appear and vice versa). Best viewed in colors.
@@
sidewalk static $16.353\%$ [Uncaptioned image] [Uncaptioned image]
@@
fence static $7.827\%$ [Uncaptioned image] [Uncaptioned image]
@@
parking static $1.681\%$ [Uncaptioned image] [Uncaptioned image]
trunk static $0.580\%$ [Uncaptioned image] [Uncaptioned image]
@@
truck dynamic $0.193\%$ [Uncaptioned image] [Uncaptioned image]
@@
motorcycle dynamic $0.045\%$ [Uncaptioned image] [Uncaptioned image]
@@
bicycle dynamic $0.018\%$ [Uncaptioned image] [Uncaptioned image]
@@
where $c$ is a small constant and $\theta$ represents the model parameters. Similar to the classic entropy minimization framework [23], the constraint in Eq. 1 is transformed into a probabilistic model where the model parameter distribution is guided by the principle of maximum entropy:
[…]
where $\lambda>0$ acts as the Lagrange multiplier associated with the entropy constraint, which corresponds to a constant $c$. $H(X_{\mathrm{in}}|A)$ has been ignored for being independent of the model parameter $\theta$. Here, we consider Eq. 2 as the formal formulation of the spatial prior and discuss how to empirically compute it in the following sections.
Marginalization. To utilize the spatial prior defined in Eq. 2, we empirically compute the entropy $H(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ of the LiDAR points inside area $A$ as follows:
[…]
where $\hat{.}$ denotes the empirical estimation. The end-to-end LiDAR segmentation model $\mathcal{G}_{\theta}$ (with parameters $\theta$) usually takes full-sized data as inputs during the inference. Therefore, to compute $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ in Eq. 3, we first pad the data outside the area to obtain the full-sized data. Here, we denote the data outside the area as $X_{\mathrm{out}}$; we then let the model infer $P(Y_{\mathrm{in}}|X_{\mathrm{in}},X_{\mathrm{out}},A)$, and finally marginalize $X_{\mathrm{out}}$ as:
@@
Training Objectives. Finally, we train the segmentation model $\mathcal{G}_{\theta}$ using the standard maximum-a-posteriori (MAP) estimation. We maximize the posterior that can be computed by Eq. 2, Eq. 3, and Eq. 4, which is formulated as follows:
@@
Here, $L(\theta)$ is the likelihood function which can be computed using labeled data, i.e., the conventional supervised learning. Minimizing $\hat{H}(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ requires the marginal probability $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ to be confident, which further requires $P(Y_{\mathrm{in}}|X_{\mathrm{in}},X_{\mathrm{out}},A)$ to be both confident and consistent with respect to different outside data $X_{\mathrm{out}}$. To sum up, our proposed semi-supervised learning framework in Eq. 5 encourages the segmentation model to make confident and consistent predictions in a predefined area, regardless of the data outside the area. The predefined area set $A$ determines the “strength” of the prior. When setting $A$ to the full area (i.e., the whole point cloud), our framework degrades to the classic entropy minimization framework [23].
@@ 3.2 LaserMix
where $(p^{x},p^{y},p^{z})$ represent the Cartesian coordinates of the LiDAR points. For any two LiDAR scans, $x_{1}$ and $x_{2}$, we first group all points from each scan by their inclination angles. More concretely, to establish $m$ non-overlapping areas, a set of $m+1$ inclination angles $\Phi=\{\phi_{0},\phi_{1},\phi_{2},...,\phi_{m}\}$ will be evenly sampled within the range of the minimum and maximum inclination angles in the dataset, and the area set $A=\{a_{1},a_{2},...,a_{m}\}$ can then be confined by bounding area $a_{i}$ in the inclination range $\phi_{i-1},\phi_{i})$. It is important to note that the range of inclination angles varies depending on the specific configurations of different LiDAR sensors.
Role in our framework: The proposed laser-based partitioning aims to effectively “excite” a strong spatial prior in the LiDAR point cloud, as described by Step 1 in our semi-supervised learning framework. As shown in Figure 1(a) and Table 1, this partitioning reveals clear distribution patterns in the semantic classes detected by each laser beam. Despite being an empirical choice, we will show in later sections that our laser-based partitioning method significantly outperforms other partition choices, including random points (MixUp-like partition [100]), random areas (CutMix-like partition [97]), and other heuristics like azimuth $\alpha$ (sensor horizontal direction) or radius $r$ (sensor range direction) partitions.
@@
[…]
where $x_{i}^{a_{j}}$ is the data crop of $x_{i}$ confined within area $a_{j}$. Correspondingly, the semantic labels are mixed in the same way as Eq. 7. It is worth highlighting that LaserMix is directly applied to the LiDAR point clouds and is thus agnostic to the various LiDAR representations, such as range view [60], bird’s eye view [102], raw points [26], sparse voxel [107], and multi-view fusion [81] representations. This versatility allows LaserMix to be effectively integrated across a broad spectrum of existing LiDAR segmentation frameworks, without necessitating any modifications to the underlying data or network structures. While LaserMix partitions the LiDAR point cloud for enhancing feature learning, concerns may arise regarding the integrity of object instances, such as cars, when these are split across different segments. However, given the large spatial area typically involved, the likelihood of significant object partitioning is low. Moreover, LaserMix’s use of both ground truth and high-confidence pseudo labels provides robust supervision, ensuring accurate object interpretation even when partitioning occurs. This design allows LaserMix to leverage the benefits of data augmentation without compromising object integrity, thereby enhancing the robustness and generalization of the model.
@@
Role in our framework: Central to our semi-supervised learning framework, LaserMix streamlines the computation of marginal probability $P(Y_{\mathrm{in}}|X_{\mathrm{in}},A)$ – a process otherwise computationally intensive in typical scenarios – as described by Step 2 in our framework. The cost for directly computing the marginal probability in Eq. 4 on real-world LiDAR data is prohibitive; we need to iterate through all areas in $A$ and all outside data in $X_{\mathrm{out}}$, which requires $|A|\cdot|X_{\mathrm{out}}|$ predictions in total. To reduce the training overhead, we take advantage of the fact that a prediction in an area will be largely affected by its neighboring areas and let $X_{\mathrm{out}}$ fill only the neighbors instead of all the remaining areas. LaserMix mixes two scans by intertwining the areas so that the neighbors of each area are filled with data from the other scan. As a result, we obtain the prediction on all areas $A$ of two scans from only two predictions, which on average reduces the cost from $|A|$ to $1$. The scan before and after mixing counts as two data fillings, therefore $|X_{\mathrm{out}}|=2$. Overall, the training overhead is reduced from $|A|\cdot|X_{\mathrm{out}}|$ to $2$: only one prediction on original data and one additional prediction on mixed data are required for each LiDAR scan. During training, the memory consumption for a batch will be $2\times$ compared to a standard semi-supervised learning framework, and the training speed will not be affected.
Figure 3: Overview of our baseline consistency regularization framework. We feed the labeled scan $x_{l}$ into the Student network to compute the supervised loss $\mathcal{L}_{\text{sup}}$ (w/ ground truth $y_{l}$). The unlabeled scan $x_{u}$ and the generated pseudo-label $y_{u}$ are mixed with $(x_{l},y_{l})$ via LaserMix (Section 3.2) to produce mixed data sample $(x_{\text{mix}},y_{\text{mix}})$, which is then fed into the Student network to compute the mixing loss $\mathcal{L}_{\text{mix}}$. Additionally, we encourage the consistency between the Student network and the Teacher network by computing the mean teacher loss $\mathcal{L}_{\text{mt}}$ over their predictions, where the Teacher network’s parameters are updated by the exponential moving average of that of the Student network. During inference, only the Teacher network is needed, which maintains the same computational cost as the conventional LiDAR segmentation pipeline.
@@
[…]
@@
Role in our framework: Our consistency regularization pipeline is designed to effectively minimize the marginal entropy for data-efficient 3D scene understanding, as described in Step 3 of our framework. Since the objective for minimizing the entropy has a challenging optimization landscape, pseudo-labeling emerges as a resort in practice [43]. Unlike conventional pseudo-label optimization in semi-supervised learning, which only aims to encourage the predictions to be confident, minimizing the marginal entropy requires all predictions to be both confident and consistent. To this end, we leverage both ground truth and pseudo-labels as anchors to encourage the 3D scene understanding model’s predictions to be confident and consistent with these supervision signals.